• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name bogdanw

    Log entry time 16:09:38 on January24,2005

    Entry number 139953

    keyword=day shift summary

    8 am, good beam
    run #2052 start 7:26 ended 8:19 250k events
    run #2053 start 8:20 ended 9:12 250k events
    run #2054 start 9:13 ended 10:20 250k events
    run #2055 start 10:21 ended 11:19 240k events
    11:12 lost beam due to south linac problem
    moved target to LD2 15 cm, raster ON 2x2 mm, 10 mkA
    run #2056 start 12:35 ended 1:46 82k events
    12:45 LH2 loop gas pressure drops, alarm ON
    page J_P, stop the beam.
    1:07 requested control access.
    Dave Meekins continues working on Loop 2 trouble, through end of shift.